OpenStudy (anonymous):

Solve the system of equations below. 5x-2y=18 3x+3y=15 ???

OpenStudy (anonymous):

online classes aha

OpenStudy (anonymous):

let's try solving it using elimination eliminate the variable y by multiplying the first equation by 3 and the second equation by 2.. then add both the new equations and tell me what you get..

OpenStudy (anonymous):

\[5x-2y=18\] \[3x+3y=15\] \[\frac{3}{3}x+\frac{3}{3}=\frac{15}{3}\] \[x+y=5\] \[x=5-y\] \[5(5-y)-2y=18\] \[25-5y-2y=18\] \[-7y=18-25\] \[-7y=-7\] \[y=1\] \[x=4\]

OpenStudy (anonymous):

5x - 2y = 18 --->(3) 3x + 3y = 15 -->(2) ----------------- 15x - 6y = 54 -- result of multiplying by 3 6x + 6y = 30 -- result of multiplying by 2 ----------------add 21x = 84 x = 4 now sub 4 in for x in either of the original equations 5x - 2y = 18 5(4) - 2y = 18 20 - 2y = 18 -2y = 18 - 20 -2y = -2 y = 1 mattpyne is correct....x = 4 and y = 1

OpenStudy (anonymous):

thank you !!!

OpenStudy (anonymous):

glad to help :)